Evaluate the following integrals: (i)Z xln(x + 1)dx [13 marks] (ii)Z sin3(lnx)cos2(lnx) x dx
1
Expert's answer
2020-06-29T15:16:04-0400

"(i) \int" xln(x+1)dx

= ln(x+1)"\int" xdx - "\int [\frac{d}{dx}(ln(x+1)) \int xdx]dx"

= "\frac {x^2ln(x+1)}{2}" - "\int \frac{x^2}{2(x+1)}" dx

= "\frac {x^2ln(x+1)}{2}" - "\int \frac{x^2-1+1}{2(x+1)}dx"

= "\frac {x^2ln(x+1)}{2}"

- "\frac{1}{2}" "[\int \frac{x^2-1}{(x+1)}dx + \int \frac{1}{(x+1)}dx]"

= "\frac {x^2ln(x+1)}{2}" - "\frac{1}{2}" [ "\int"(x-1)dx+ "\int\frac{1}{(x+1)}dx" ]

= "\frac {x^2ln(x+1)}{2}" - "\frac{1}{2}" [ "\frac{x^2}{2}" - x + ln(x+1)] + C

= "\frac {x^2ln(x+1)}{2}" - "\frac{1}{2}" "\frac{x^2-2x}{2}" - "\frac {ln(x+1)}{2}" + C

= "\frac {(x^2-1)ln(x+1)}{2} - \frac{x^2-2x}{4} +" C where C is integration constant.

(ii)

I = "\int" sin 3(lnx) cos 2(lnx) x dx

So 2I = "\int" 2 sin 3(lnx) cos 2(lnx) x dx

= "\int" x[sin 5(lnx) + sin (lnx)]dx

= "\int" xsin 5(lnx)dx + "\int"xsin (lnx)dx

= I5 + I1 (let)

Where I5 = "\int" xsin 5(lnx)dx

and I1 = "\int" xsin (lnx)dx

Now let us find a general formula for "I_m=" "\int" xsin m(lnx)dx

"I_m=\int" xsin m(lnx)dx

"I_m" = sin m(lnx) "\int" xdx "-"

"\int" "[\frac{d}{dx}(sin m(lnx))\int xdx]dx"

"I_m" = "\frac{x^2}{2}sin m(lnx)-\int \frac{mcos m(lnx)}{x} \frac {x^2}{2}dx"

"I_m" = "\frac{x^2}{2}sin m(lnx)-" "\frac{m}{2}\int xcos m(lnx) dx"

"I_m" = "\frac{x^2}{2}sin m(lnx)-" "\frac {m}{2} [cos m(lnx)\int xdx" - "\int (\frac{d}{dx}(cos m(lnx))\int xdx)]"

"I_m" = "\frac{x^2}{2}sin m(lnx) -" "\frac{m}{2}" ["\frac{x^2}{2}cos m(lnx)-" "\int - \frac{msin m(lnx)}{x} \frac {x^2}{2}dx]"

"I_m" = "\frac{x^2}{2}sin m(lnx)" - "\frac{m}{2}" ["\frac{x^2}{2}cos m(lnx)" + "\frac{m}{2}" "\int" x sin m(lnx)dx]

"I_m" = "\frac{x^2}{2}sin m(lnx)" - "\frac{m}{2}" "\frac{x^2}{2}cos m(lnx)" - "\frac{m^2}{4}" "I_m"

So "(1+\frac{m^2}{4})" "I_m" = "\frac{x^2}{4}[2sin m(lnx)-mcos m(lnx)]"

So "I_m" = "\frac{x^2}{m^2+4}[2sin m(lnx)-mcos m(lnx)]"

So putting m = 5 we get

I5= "\frac{x^2}{5^2+4}[2sin 5(lnx)-5cos 5(lnx)]"

= "\frac{x^2}{29}[2sin 5(lnx)-5cos 5(lnx)]"

and putting m = 1 we get

I1= "\frac{x^2}{1^2+4}[2sin (lnx)-cos (lnx)]"

= "\frac{x^2}{5}[2sin (lnx)-cos (lnx)]"

So 2I = "\frac{x^2}{29}[2sin 5(lnx)-5cos 5(lnx)]" + "\frac{x^2}{5}[2sin (lnx)-cos (lnx)]"

=> I = "\frac{x^2}{58}[2sin 5(lnx)-5cos 5(lnx)]" + "\frac{x^2}{10}[2sin (lnx)-cos (lnx)]" + C

So "\int" sin 3(lnx) cos 2(lnx) x dx

= "\frac{x^2}{58}[2sin 5(lnx)-5cos 5(lnx)]" + "\frac{x^2}{10}[2sin (lnx)-cos (lnx)]" + C

where C is integration constant